OpenStudy (anonymous):

Find the values of p for which the equation (p+1)x^2 + 4px +9 = 0 has equal roots

OpenStudy (ash2326):

ax^2+bx+c=0 for equal roots D=0 i.e b^2-4ac=0 here b= 4p a=p+1 c=9 so (4p)^2-4(p+1)9=0 16p^2-36p-36=0 divide whole equation by4 4p^2-9p-9=0 4p^2-12p+3p-9=0 4p(p-3)+3(p-3)=0 (4p+3)(p-3)=0 p=-3/4 or p=3

OpenStudy (anonymous):

thanks

OpenStudy (ash2326):

welcome

OpenStudy (anonymous):

Also, (b) Find the range of values of m for which the straight line y=mx-5 intesects the curve y=x^2-1 in two disctint points?

OpenStudy (ash2326):

substitute y=mx-5 in the second equation, in last question we wanted only one root , here we need two distinct roots for that we need to have D>=0 mx-5=x^2-1 x^2-mx+4=0 ax^2+bx+c=0 D>0 b62-4ac>0 m^2-4*1*4>0 m^2-16>0 (m-4)(m+4)>0 either m<-4 or m.>4

OpenStudy (ash2326):

wrote D>=0 it should be D>0, sorry :-|

OpenStudy (ash2326):